OpenStudy (anonymous):

What are the possible values of x in x^2+3x+3=0?

OpenStudy (campbell_st):

so you are being asked to solve the quadratic equation... and you will need the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 1, b = 3 and c = 3 you need to be aware that the solutions are complex numbers